Show that the general solution of the differential equation

4z'' - 4z' - 3z = cos2x takes the expression: z = c1e^(3x/2) + c2e^(-x/2) - (19/425)cos2x - (8/425)sin2x


1
Expert's answer
2021-03-26T08:38:16-0400

Solution.

"4z''-4z'-3z=\cos{2x}."


This is a linear differential equation of the second-order with constant coefficients. The general solution of this equation is found as the sum of the general solution "\tilde{z}" ​ of the corresponding homogeneous equation and some particular integral solution "z_*\u200b" of the inhomogeneous equation:"zy=\tilde{z}+z_*."

Сompose and solve the characteristic equation:



"4\lambda^2-4\lambda-3=0, \newline
\lambda_1=-\frac{1}{2}, \lambda_2=\frac{3}{2}."

 Therefore, "\tilde{z}=C_1e^{\lambda_1x}+C_2e^{\lambda_2x}." So,


"\tilde{z}=C_1e^{-\frac{1}{2}x}+C_2e^{\frac{3}{2}x}."

Since the right-hand side contains a "\cos{2x}" , we find a particular integral solution "z_*" ​ ​in the form: 


"z_*=A\cos{2x}+B\sin{2x}."


Using the method of undetermined coefficients we find that "A=-\frac{19}{425}, B=-\frac{8}{425}."

So, the particular integral solution of equation "4z''-4z'-3z=\cos{2x}" is


"z_*=-\frac{19}{425}\cos{2x}-\frac{8}{425}\sin{2x}."


And the general solution of the same equation is


"z=C_1e^{-\frac{1}{2}x}+C_2e^{\frac{3}{2}x}-\frac{19}{425}\cos{2x}-\frac{8}{425}\sin{2x}."

Answer. "z=C_1e^{-\frac{1}{2}x}+C_2e^{\frac{3}{2}x}-\frac{19}{425}\cos{2x}-\frac{8}{425}\sin{2x}."